# Semi-Supervised Clustering of Sparse Graphs: Crossing the Information-Theoretic Threshold

### 08 Jan 2024 v3

@@ 1.4 Semi-Supervised Learning
Within machine learning there are three basic approaches: supervised learning, unsupervised learning and the combination of both, semi-supervised learning. The main difference lies in the availability of labeled data. While unsupervised learning (e.g. clustering, association and dimension reduction) operates without any domain-specific guidance or preexisting knowledge, supervised learning (e.g. classification and regression) relies on all training samples being associated with labels. However, it is often the case where existing knowledge for a problem domain doesn’t fit either of these extremes.
@@
In real-world applications, unlabeled data comes with a much lower cost not requiring expensive human annotation and laboratory experiments. For example, documents crawled from the Web, images obtained from surveillance cameras, and speech collected from broadcast are relatively more accessible comparing to their labels which are required for prediction tasks, such as sentiment orientation, intrusion detection, and phonetic transcript. Motivated by this labeling bottleneck, the semi-supervised approach utilizes both labeled and unlabeled data to perform learning tasks faster, better, and cheaper. Since the 1990s, semi-supervised learning research has enjoyed an explosion of interest with applications like natural language processing (Qiu et al., 2019; Chen and Xu, 2016) and computer vision (Xie et al., 2020; Lee, 2013).
@@
This paper is closely related to the subtopic called constrained clustering, where one has some must-links (i.e. two nodes belong to the same cluster) and cannot-links (i.e. two nodes are in different clusters) as extra information. Although constrained versions of classic algorithms have been studied empirically, such as expectation–maximization (Shental et al., 2003), k-means (Wagstaff et al., 2001) and spectral method (Kamvar et al., 2003), our methods take different approaches than hard-coding these pairwise constraints into the algorithms and provide theoretically insightful guarantees under SBM.
@@
There also has been some excellent work considering introducing node information into SBMs. One interesting direction is overlaying a Gaussian mixture model with SBM, namely at each node of the graph assuming there is a vector of Gaussian covariates, which are correlated with the community structure. (Yan and Sarkar, 2016) proposed an SDP with k-means regularization and showed that such node information indeed improves the clustering accuracy, while (Lu and Sen, 2020) formally established the information theoretic threshold for this model when the average degree exceeds one. The node information is depicted as noisy observations for all nodes in (Mossel and Xu, 2015). In this setting, random guessing is obviously no longer a meaningful baseline. Hence, the authors refer to the Maximum A Posteriori (MAP) estimator instead, and show a local belief propagation algorithm’s performance converges to the MAP accuracy in various regimes of the model. They also conjectured that those regimes can be extended to the entire domain of $a>b$ with arbitrary, but non-vanishing, strength of node information. We will see in the next section that our result establishes this conjecture in the sense that with arbitrary, but non-vanishing, knowledge of the labels, we can beat the meaningful baseline for all $a>b$.
@@
All of these models require input on every node, which does not really fall within the scope of semi-supervised learning. Whereas, we consider a realistic and intuitive generalization of SBM where a small random fraction of the labels is given accurately. (Kanade et al., 2014) studied basically the same model as ours. They demonstrated that a vanishing fraction of labels improves the power of local algorithms. In particular, when the number of clusters diverges, it helps their local algorithm to go below the conjectured algorithmic threshold of the unsupervised case. Elegantly, they also proved the following result, which is closely related to this paper.
@@
(Kanade et al., 2014) When the fraction of revealed node labels is vanishingly small, (unsupervised) weak recovery problem on the planted bisection model is not solvable if SNR is under the Kesten-Stigum threshold.
@@ 2 Our Results
The main goal of this paper is to answer the long-standing open question regarding the semi-supervised learning on probabilistic graph models. We would like to quote the version from (Abbe, 2018):
@@
In the previous section, we have discussed deep research related to the clustering / community detection problem on the SBM. Establishing the phase transition phenomenon at KS threshold is a major focal point. However, such a sharp and intrinsic limit totally disappears when an arbitrarily small fraction of the labels is revealed. This astonishing change is first observed in (Zhang et al., 2014) where the authors provide non-trivial conjectures based on calculations of the belief propagation approach.
The theory of semi-supervised clustering contains some fascinating and fundamental algorithmic challenges arising from both the sparse random graph model itself and the semi-supervised learning perspective. To address them rigorously, we first define the semi-supervised SBM in a way that it captures the essence of realistic semi-supervised learning scenarios and is a natural and simple generalization of unsupervised models.
@@
For $n\in\mathbb{N}$, $p,q\in(0,1)$ and $\rho\geq 0$, let $\operatorname{\mathcal{G}}(n,p,q,\rho)$ denote the distribution over graphs with $n$ vertices and $n$-dimensional vectors defined as follows. The vertex set is partitioned uniformly at random into two subsets $S_{1},S_{2}$ under the balance constraint $|S_{1}|=|S_{2}|=n/2$. Then conditional on the partition, two processes are undertaken independently:
  • Let $E$ denote the edge set of the graph $G$. Edges are included independently with probability defined as follows.
@@
    $\tilde{x}_{i}=\begin{cases}~{}1,\quad&i\in\mathcal{R}\cap S_{1},\\ -1,\quad&i\in\mathcal{R}\cap S_{2},\\ ~{}0,\quad&\text{otherwise.}\end{cases}$ (9)
@@
The revealing process is independent from edge formation, i.e., $G\perp\tilde{x}|S_{1},S_{2}$. Moreover, if we set $\rho=0$ or simply ignore the revealed labels, the model is exactly the unsupervised SBM. In other words, the marginal distribution of the random graph is indeed $\operatorname{\mathcal{G}}(n,p,q)$ from Definition 1.
@@
The definition is versatile in the sense that it keeps unsupervised setting as a special case (and with it all the interesting phase transitions). On the other hand, it can be easily generalized to the multiple and/or asymmetric communities case.
Under semi-supervised setting, we naturally extend community detection problems in a non-trivial way that includes the unsupervised situation as a special case and captures its essence. We will discuss these items in detail when it comes to the corresponding section.
@@
Semi-supervised distinguishability: finding a test that with high probability, tells $\operatorname{\mathcal{G}}(n,d/n,d/n,\boldsymbol{\rho})$ from $\operatorname{\mathcal{G}}(n,a/n,b/n,\boldsymbol{\rho})$ where $d=\frac{a+b}{2},~{}a>b$.
@@
$\hat{x}_{v}=\text{sgn}\left(\sum_{i\in\{u\in\mathcal{R}\mathrel{\mathop{:}}~{}d(u,v)=t\}}x_{i}\right)$ (10)
@@
Although it successfully solves the weak recovery problem in the largest possible regime, there are some limitations hindering the census method’s utility in practice. Its performance depends on a sufficient amount of revealed labels, hence requiring $n$ to be really large. Besides, without an unsupervised counterpart, it is not applicable when revealing is unreliable.
To address these challenges, we propose our second semi-supervised clustering algorithm which performs well in practice and covers the unsupervised setting as special case. As discussed in the previous section, SDPs enjoy many nice properties, among which the monotone-robustness is particularly interesting to us. In the semi-supervised setting, the revealed labels are supposed to enhance the community structure. However, the work from (Moitra et al., 2016) suggests such enhancement may not help with, but to the contrary can hurt the performance of many algorithms, which makes SDP an ideal starting point for us. We define the Constrained Semidefinite Program (CSDP) as
@@ 2.1 Proof Techniques
The technical challenges of establishing Theorem 11 are mainly due to the fact that the advantage created by revealed labels can be easily blurred out by various approximations of the limit distribution. Instead of the central limit theorem, one needs a Berry–Esseen-type inequality to derive a more quantitative result of the convergence rate. Moreover, since the distribution of each underlying component also depends on $n$, the conventional sample mean formulation does not apply here. We overcome the difficulty above with a direct analysis of non-asymptotic distributions, which leads to a detailed comparison between two binomial variable sequences with constant expectations.
It is quite surprising that this calculation can be carried out in rather elegant manner, since many other applications of this method are much more technically involved. For example to establish the independence among estimators, one may need to consider the ‘leave-one-out’ trick. But in our case, it comes in a very natural way.
Regarding CSDP, we first show it can be coupled to a SDP with the surrogate input matrices. Moreover, its optimal value lies in between two unsupervised SDPs associated with the same random graph model (different parameters). This means all the analytical results from SDP research can be transferred into the CSDP study. However, we notice that it is common to make assumptions on the average degree $d$ in the relevant literature. It is quite reasonable in the unsupervised setting since the graph topology is a strong indicator for the possibility of weak recovery, e.g. when $d\leq 1$, there will not exist a giant component that is of size linear in $n$.
To establish our result without such extra assumptions, we derive a probabilistic bound on the cut norm of centered adjacency matrix and then use Grothendieck’s inequality to bound the SDP on ERMs from above. This idea follows from (Guédon and Vershynin, 2014); we give a slightly different analysis fitting for our purpose. A generalized weak law of large number is also derived to address the issue that distributions of the entries change as $n\to\infty$. Then we conclude the proof with a lower bound of the CSDP on SBMs considering a witness that consists of the ground truth of labels.
@@ 2.3 Notation
For any $n\in\mathbb{N}$, we denote the first $n$ integers by $[n]=\{1,2,\dots,n\}$. For a set $S$, its cardinality is denoted by $|S|$. We use lowercase letters for vectors (e.g. $v=(v_{1},v_{2},\dots,v_{n})$) and uppercase letters for matrices (e.g. $M=[M_{ij}]_{i,j\in[n]}$). In particular, for adjacency matrices, we omit their dependency on underlying graphs. Instead of $A_{G}$, we simply write $A$. $\mathbf{1}_{n}=(1,\dots,1)\in\mathbb{R}^{n}$ stands for the all-ones vector and $I_{n}$ is the $n\times n$ identity matrix. $\mathbf{e}_{i}\in\mathbb{R}_{n}$ represents the $i$’s standard basis vector. For two real-valued matrices $A$ and $B$ with the same dimensions, we define the Frobenius inner product as $\langle A,B\rangle=\sum_{i,j}A_{ij}\cdot B_{ij}=\operatorname{Tr}(A^{\top}B)$. Vector inner product is viewed as a special case of $n\times 1$ matrices. Let $\|v\|_{p}=(\sum_{i=1}^{p}\|v_{i}\|^{p})^{1/p}$ be the $\ell_{p}$ norm of vectors with standard extension to $p=\infty$. Let $\|M\|_{p\to q}=\sup_{\|v\|p\leq 1}\|Mv\|_{q}$ be the $\ell_{p}$-to-$\ell_{q}$ operator norm and $\|M\|_{\text{op}}\coloneqq\|M\|_{2}\coloneqq\|M\|_{2\to 2}$. Random graphs induce measures on the product space of label, edge and revealed node assignments over $n$ vertices. For any $n\in\mathbb{N}$, it is implicitly understood that one such measure is specified with that graph size. The terminology with high probability means ‘with probability converging to $1$ as $n\to\infty$’. Also, we follow the conventional Big-Oh notation for asymptotic analysis. $o_{p}(1)$ stands for convergence to 00 in probability.
@@ 3 Census Method
Analysis of model from Definition 1 is a challenging task since conditioned on the graph, it is neither an Ising model, nor a Markov random field. This is mainly due to following facts: (1) The balance requirement puts a global condition on the size of each cluster; (2) Even if conditioned on sizes, there is a slight repulsion between unconnected nodes. Namely, if two nodes do not form an edge, the probability of them being in the same community is different from them being in the opposite communities.
Recent years have witnessed a series of excellent contributions on the phase transitions in the sparse regime. Our census method for semi-supervised clustering is mainly inspired by the natural connection between community detection on SBMs and reconstruction on trees, which is formally established in (Mossel et al., 2015). Intuitively, for a vertex $v$ in $\operatorname{\mathcal{G}}(n,a/n,b/n)$, it is not likely that a node from its small neighborhood has an edge leading back to $v$. Therefore, the neighborhood looks like a random labelled tree with high probability. Furthermore, the labelling on the vertices behaves like the broadcasting a bit from the root of a tree down to its leaves (see the survey (Mossel, 2001) for a detailed discussion).
@@ 3.1 Majority of t-Neighbors
Let $(G,x)$ obey the planted bisection model $\operatorname{\mathcal{G}}(n,a/n,b/n)$. We denote the set of all vertices by $V(G)$. For a fixed vertex $v$ and $t\in\mathbb{N}$, let $N_{t}(v)$ denote the number of vertices which are $t$ edges away from $v$. $\Delta_{t}(v)$ is defined as the difference between the numbers of $t$-neighbors in each community. Namely,
@@
If one assumes that the subgraph of $G$ induced by the vertices within $t$ edges of $v$ is a tree, the expected value of $N_{t}(v)$ is approximately $[(a+b)/2]^{t}$ and the expected value of $x_{v}\cdot\Delta_{t}(v)$, i.e., the expected number of these vertices in the same community as $v$ minus the expected number of these vertices in the other community, is approximately $[(a-b)/2]^{t}$. So, if one can somehow independently determine which community a vertex is in with an accuracy of $1/2+\alpha$ for some $\alpha>0$, one will be able to predict the label of each vertex with an accuracy of roughly $1/2+[(a-b)^{2}/(2(a+b))]^{t/2}\cdot\alpha$, by guessing it as the majority of $v$’s $t$-neighbors. Under the unsupervised learning setting, one can get a small advantage, $\alpha\sim\Theta(1/\sqrt{n})$, by randomly initializing labels. It is guaranteed by the central limit theorem that such a fraction exists in either an agreement or disagreement form.
Figure 2: Neighborhood of node $v$ with a tree structure. True clusters are coded in black and white. The shaded area indicates those nodes randomly guessed to be in the same community or the opposite community as $v$. The annulus represents the collection of its $t$-neighbors.
To amplify this lucky guess, we need $t$ to be sufficiently large so that $[(a-b)^{2}/(2(a+b))]^{t/2}>\sqrt{n}$, which implies $[(a+b)/2]^{t}>n$. Note $d=(a+b)/2$ is the average degree. This means before the signal is strong enough for our purpose, not only our tree approximation will break down, but vertices will be exhausted. However, if we have access to some of the true labels, i.e., in the semi-supervised setting, we can leverage the tree structure for a non-vanishing advantage over random guessing.
@@
$Y_{u}=\begin{cases}x_{u}&\text{if~{}}A_{uv}=1\\ 0&\text{otherwise}\end{cases}$ (14)
@@
$\operatorname{P}(Y_{u}=1|x_{v}=1)=\frac{\operatorname{P}(Y_{u}=1,x_{v}=1)}{\operatorname{P}(x_{v}=1)}\approx\frac{a}{2n}$ (17)
@@
$\operatorname{P}(Y_{u}=-1|x_{v}=1)\approx\frac{b}{2n}$ (18)
@@
$\tilde{\Delta}_{1}(v)=\sum_{u\in\mathcal{R}}Y_{u}$ (20)
@@
$\operatorname{E}(\tilde{\Delta}_{1}(v))=\sum_{u\in\mathcal{R}}\operatorname{E}(Y_{u})=\rho\frac{a-b}{2}$ (21)
$\operatorname{Var}(\tilde{\Delta}_{1}(v))=\sum_{u\in\mathcal{R}}\operatorname{Var}(Y_{u})=\rho\frac{a+b}{2}+o(1)$ (22)

### 28 Feb 2024 v4

@@ 1.4 Semi-Supervised Learning
Within machine learning, there are three basic approaches: supervised learning, unsupervised learning, and the combination of both, semi-supervised learning. The main difference lies in the availability of labeled data. While unsupervised learning (e.g., clustering, association, and dimension reduction) operates without any domain-specific guidance or preexisting knowledge, supervised learning (e.g., classification and regression) relies on all training samples being associated with labels. However, it is often the case where existing knowledge for a problem domain doesn’t fit either of these extremes.
@@
In real-world applications, unlabeled data comes with a much lower cost not requiring expensive human annotation and laboratory experiments. For example, documents crawled from the Web, images obtained from surveillance cameras, and speech collected from the broadcast are relatively more accessible compared to their labels which are required for prediction tasks, such as sentiment orientation, intrusion detection, and phonetic transcript. Motivated by this labeling bottleneck, the semi-supervised approach utilizes both labeled and unlabeled data to perform learning tasks faster, better, and cheaper. Since the 1990s, semi-supervised learning research has enjoyed an explosion of interest with applications like natural language processing (Qiu et al., 2019; Chen and Xu, 2016) and computer vision (Xie et al., 2020; Lee, 2013).
@@
This paper is closely related to the subtopic called constrained clustering, where one has some must-links (i.e., two nodes belong to the same cluster) and cannot-links (i.e., two nodes are in different clusters) as extra information. Although constrained versions of classic algorithms have been studied empirically, such as expectation–maximization (Shental et al., 2003), k-means (Wagstaff et al., 2001), and spectral method (Kamvar et al., 2003), our methods take different approaches than hard-coding these pairwise constraints into the algorithms and provide theoretically insightful guarantees under SBM.
@@
There also has been some excellent work considering introducing node information into SBMs. One interesting direction is overlaying a Gaussian mixture model with SBM, namely at each node of the graph assuming there is a vector of Gaussian covariates, which are correlated with the community structure. (Yan and Sarkar, 2016) proposed an SDP with k-means regularization and showed that such node information indeed improves the clustering accuracy, while (Lu and Sen, 2020) formally established the information-theoretic threshold for this model when the average degree exceeds one. The node information is depicted as noisy observations for all nodes in (Mossel and Xu, 2015). In this setting, random guessing is no longer a meaningful baseline. Hence, the authors refer to the Maximum A Posteriori (MAP) estimator instead and show a local belief propagation algorithm’s performance converges to the MAP accuracy in various regimes of the model. They also conjectured that those regimes can be extended to the entire domain of $a>b$ with arbitrary, but non-vanishing, strength of node information. We will see in the next section that our result establishes this conjecture in the sense that with arbitrary, but non-vanishing, knowledge of the labels, we can beat the meaningful baseline for all $a>b$.
@@
All of these models require input on every node, which does not fall within the scope of semi-supervised learning. Whereas, we consider a realistic and intuitive generalization of SBM where a small random fraction of the labels is given accurately. (Kanade et al., 2014) studied the same model as ours. They demonstrated that a vanishing fraction of labels improves the power of local algorithms. In particular, when the number of clusters diverges, it helps their local algorithm to go below the conjectured algorithmic threshold of the unsupervised case. Elegantly, they also proved the following result, which is closely related to this paper.
@@
(Kanade et al., 2014) When the fraction of revealed node labels is vanishingly small, the (unsupervised) weak recovery problem on the planted bisection model is not solvable if SNR is under the Kesten-Stigum threshold.
@@ 2 Our Results
The main goal of this paper is to answer the long-standing open question regarding semi-supervised learning on probabilistic graph models. We would like to quote the version from (Abbe, 2018):
@@
In the previous section, we discussed deep research related to the clustering/community detection problem on the SBM. Establishing the phase transition phenomenon at the KS threshold is a major focal point. However, such a sharp and intrinsic limit completely disappears when an arbitrarily small fraction of the labels is revealed. This astonishing change is first observed in (Zhang et al., 2014) where the authors provide non-trivial conjectures based on calculations of the belief propagation approach.
The theory of semi-supervised clustering contains some fascinating and fundamental algorithmic challenges arising from both the sparse random graph model itself and the semi-supervised learning perspective. To address them rigorously, we first define the semi-supervised SBM so that it captures the essence of realistic semi-supervised learning scenarios and is a natural and simple generalization of unsupervised models.
@@
For $n\in\mathbb{N}$, $p,q\in(0,1)$ and $\rho\geq 0$, let $\operatorname{\mathcal{G}}(n,p,q,\rho)$ denote the distribution over graphs with $n$ vertices and $n$-dimensional vectors defined as follows. The vertex set is partitioned uniformly at random into two subsets $S_{1},S_{2}$ under the balance constraint $|S_{1}|=|S_{2}|=n/2$. Then, conditioned on the partition, two processes are undertaken independently:
  • Let $E$ denote the edge set of the graph $G$. Edges are included independently with probability defined as follows:
@@
    $\tilde{x}_{i}=\begin{cases}~{}1\quad&i\in\mathcal{R}\cap S_{1},\\ -1\quad&i\in\mathcal{R}\cap S_{2},\\ ~{}0\quad&\text{otherwise.}\end{cases}$ (9)
@@
The revealing process is independent of edge formation, i.e., $G\perp\tilde{x}|S_{1},S_{2}$. Moreover, if we set $\rho=0$ or simply ignore the revealed labels, the model is exactly the unsupervised SBM. In other words, the marginal distribution of the random graph is indeed $\operatorname{\mathcal{G}}(n,p,q)$ from Definition 1.
@@
The definition is versatile in the sense that it keeps the unsupervised setting as a special case (and with it all the interesting phase transitions). On the other hand, it can be easily generalized to the multiple and/or asymmetric communities case.
Under the semi-supervised setting, we naturally extend community detection problems in a non-trivial way that includes the unsupervised situation as a special case and captures its essence. We will discuss these items in detail when it comes to the corresponding section.
@@
Semi-supervised distinguishability: finding a test that, with high probability, distinguishes $\operatorname{\mathcal{G}}(n,d/n,d/n,\boldsymbol{\rho})$ from $\operatorname{\mathcal{G}}(n,a/n,b/n,\boldsymbol{\rho})$ where $d=\frac{a+b}{2},~{}a>b$.
@@
$\hat{x}_{v}=\text{sgn}\left(\sum_{i\in\{u\in\mathcal{R}\mathrel{\mathop{:}}~{}d(u,v)=t\}}x_{i}\right),$ (10)
@@
Furthermore, we derive an explicit constant bound for the overlap with the corresponding tail probability. Namely, under $\operatorname{\mathcal{G}}(n,a/n,b/n,\rho)$, if we denote Overlap on the unrevealed nodes, between 1-neighbors census estimator and the ground truth, as $\Theta$ and let $\delta=\frac{\rho(a-b)}{2\mathrm{e}^{\rho(a+b)}}$, then $\operatorname{P}\left(\Theta\geq\frac{\delta}{2}\right)\geq 1-\mathrm{e}^{-\frac{\delta^{2}(1-\rho)n}{8}}$. A detailed discussion is provided in the next section.
@@
Although it successfully solves the weak recovery problem in the largest possible regime, some limitations are hindering the census method’s utility in practice. Its performance depends on a sufficient amount of revealed labels, hence requiring $n$ to be quite large. Besides, without an unsupervised counterpart, it is not applicable when revealing is unreliable.
To address these challenges, we propose our second semi-supervised clustering algorithm which performs well in practice and covers the unsupervised setting as a special case. As discussed in the previous section, SDPs enjoy many nice properties, among which the monotone-robustness is particularly interesting to us. In the semi-supervised setting, the revealed labels are supposed to enhance the community structure. However, the work from (Moitra et al., 2016) suggests such enhancement may not help with, but to the contrary can hurt the performance of many algorithms, which makes SDP an ideal starting point for us. We define the Constrained Semidefinite Program (CSDP) as
@@ 2.1 Proof Techniques
The technical challenges of establishing Theorem 11 root in the fact that the advantage created by revealed labels can be easily blurred out by various approximations of the limit distribution. Instead of the central limit theorem, one needs a Berry–Esseen-type inequality to derive a more quantitative result of the convergence rate. Moreover, since the distribution of each underlying component also depends on $n$, the conventional sample mean formulation does not apply here. We overcome the difficulty above with a direct analysis of non-asymptotic distributions, which leads to a detailed comparison between two binomial variable sequences with constant expectations.
It is quite surprising that this calculation can be carried out in a rather elegant manner since many other applications of this method are much more technically involved. For example, to establish independence among estimators, one may need to consider the ‘leave-one-out’ trick. But in our case, it comes in a very natural way.
Regarding CSDP, we first show it can be coupled to an SDP with the surrogate input matrices. Moreover, its optimal value lies between two unsupervised SDPs associated with the same random graph model (different parameters). It means that all the analytical results from SDP research can be transferred into the CSDP study. However, we notice that it is common to make assumptions on the average degree $d$ in the relevant literature. It is quite reasonable in the unsupervised setting since the graph topology is a strong indicator of the possibility of weak recovery. E.g., when $d\leq 1$, there will not exist a giant component that is of size linear in $n$.
To establish our result without such extra assumptions, we derive a probabilistic bound on the cut norm of the centered adjacency matrix and then use Grothendieck’s inequality to bound the SDP on ERMs from above. This idea follows from (Guédon and Vershynin, 2014); we give a slightly different analysis fitting for our purpose. A generalized weak law of large numbers is also derived to address the issue that distributions of the entries change as $n\to\infty$. Then we conclude the proof with a lower bound of the CSDP on SBMs considering a witness that consists of the ground truth of labels.
@@ 2.3 Notation
For any $n\in\mathbb{N}$, we denote the first $n$ integers by $[n]=\{1,2,\dots,n\}$. For a set $S$, its cardinality is denoted by $|S|$. We use lowercase letters for vectors (e.g., $v=(v_{1},v_{2},\dots,v_{n})$) and uppercase letters for matrices (e.g., $M=[M_{ij}]_{i,j\in[n]}$). In particular, for adjacency matrices, we omit their dependency on underlying graphs. Instead of $A_{G}$, we simply write $A$. $\mathbf{1}_{n}=(1,\dots,1)\in\mathbb{R}^{n}$ stands for the all-ones vector, and $I_{n}$ is the $n\times n$ identity matrix. $\mathbf{e}_{i}\in\mathbb{R}_{n}$ represents the $i$’s standard basis vector. For two real-valued matrices $A$ and $B$ with the same dimensions, we define the Frobenius inner product as $\langle A,B\rangle=\sum_{i,j}A_{ij}\cdot B_{ij}=\operatorname{Tr}(A^{\top}B)$. Vector inner product is viewed as a special case of $n\times 1$ matrices. Let $\|v\|_{p}=(\sum_{i=1}^{p}\|v_{i}\|^{p})^{1/p}$ be the $\ell_{p}$ norm of vectors with standard extension to $p=\infty$. Let $\|M\|_{p\to q}=\sup_{\|v\|p\leq 1}\|Mv\|_{q}$ be the $\ell_{p}$-to-$\ell_{q}$ operator norm and $\|M\|_{\text{op}}\coloneqq\|M\|_{2}\coloneqq\|M\|_{2\to 2}$. Random graphs induce measures on the product space of label, edge, and revealed node assignments over $n$ vertices. For any $n\in\mathbb{N}$, it is implicitly understood that one such measure is specified with that graph size. The terminology with high probability means ‘with probability converging to $1$ as $n\to\infty$’. Also, we follow the conventional Big-Oh notation for asymptotic analysis. $o_{p}(1)$ stands for convergence to 00 in probability.
@@ 3 Census Method
Analysis of the model from Definition 1 is a challenging task since conditioned on the graph, it is neither an Ising model nor a Markov random field. This is mainly due to the following facts: (1) The balance requirement puts a global condition on the size of each cluster; (2) Even if conditioned on sizes, there is a slight repulsion between unconnected nodes. Namely, if two nodes do not form an edge, the probability of them being in the same community is different from the probability of them being in opposite communities.
Recent years have witnessed a series of excellent contributions to the study of phase transitions in the sparse regime. Our census method for semi-supervised clustering is mainly inspired by the natural connection between community detection on SBMs and reconstruction problems on trees, which was formally established by (Mossel et al., 2015). Intuitively, for a vertex $v$ in $\operatorname{\mathcal{G}}(n,a/n,b/n)$, it is not likely that a node from its small neighborhood has an edge leading back to $v$. Therefore, the neighborhood looks like a randomly labeled tree with high probability. Furthermore, the labeling on the vertices behaves like broadcasting a bit from the root of a tree down to its leaves (see the survey (Mossel, 2001) for a detailed discussion).
@@ 3.1 Majority of t-Neighbors
Let $(G,x)$ obey the planted bisection model $\operatorname{\mathcal{G}}(n,a/n,b/n)$. We denote the set of all vertices by $V(G)$. For a fixed vertex $v$ and $t\in\mathbb{N}$, let $N_{t}(v)$ denote the number of vertices that are $t$ edges away from $v$. $\Delta_{t}(v)$ is defined as the difference between the numbers of $t$-neighbors in each community. Namely,
@@
If one assumes that the subgraph of $G$ induced by the vertices within $t$ edges of $v$ is a tree, the expected value of $N_{t}(v)$ is approximately $[(a+b)/2]^{t}$, and the expected value of $x_{v}\cdot\Delta_{t}(v)$, i.e., the expected number of these vertices in the same community as $v$ minus the expected number of these vertices in the other community, is approximately $[(a-b)/2]^{t}$. So, if one can somehow independently determine which community a vertex is in with an accuracy of $1/2+\alpha$ for some $\alpha>0$, one will be able to predict the label of each vertex with an accuracy of roughly $1/2+[(a-b)^{2}/(2(a+b))]^{t/2}\cdot\alpha$, by guessing it as the majority of $v$’s $t$-neighbors. Under the unsupervised learning setting, one can get a small advantage, $\alpha\sim\Theta(1/\sqrt{n})$, by randomly initializing labels. It is guaranteed by the central limit theorem that such a fraction exists in either an agreement or disagreement form.
Figure 2: Neighborhood of node $v$ with a tree structure. The ground truth of clusters is coded in black and white. The shaded area indicates those nodes randomly guessed to be in the same community or the opposite community as $v$. The annulus represents the collection of its $t$-neighbors.
To amplify this lucky guess, we need $t$ to be sufficiently large so that $[(a-b)^{2}/(2(a+b))]^{t/2}>\sqrt{n}$, which implies $[(a+b)/2]^{t}>n$. Note that $d=(a+b)/2$ is the average degree. This means before the signal is strong enough for our purpose, not only our tree approximation will break down, but vertices will also be exhausted. However, if we have access to some of the true labels, i.e., in the semi-supervised setting, we can leverage the tree structure to get a non-vanishing advantage over random guessing.
@@
$Y_{u}=\begin{cases}x_{u}&\text{if~{}}A_{uv}=1,\\ 0&\text{otherwise}.\end{cases}$ (14)
@@
$\operatorname{P}(Y_{u}=1|x_{v}=1)=\frac{\operatorname{P}(Y_{u}=1,x_{v}=1)}{\operatorname{P}(x_{v}=1)}\approx\frac{a}{2n}.$ (17)
@@
$\operatorname{P}(Y_{u}=-1|x_{v}=1)\approx\frac{b}{2n}.$ (18)
@@
$\tilde{\Delta}_{1}(v)=\sum_{u\in\mathcal{R}}Y_{u}.$ (20)
@@
$\operatorname{E}(\tilde{\Delta}_{1}(v))=\sum_{u\in\mathcal{R}}\operatorname{E}(Y_{u})=\rho\frac{a-b}{2},$ (21)
$\operatorname{Var}(\tilde{\Delta}_{1}(v))=\sum_{u\in\mathcal{R}}\operatorname{Var}(Y_{u})=\rho\frac{a+b}{2}+o(1).$ (22)